# Practitioner Motives to Use Different Hyperparameter Optimization Methods

### v3

@@ 4.2. Perceived Success of Using HPO Methods to Achieve Specific Goals
Practitioners appear to have individual motives to use specific HPO methods. Yet, some decisions for HPO methods may not produce the attempted results. To distinguish between successful and unsuccessful experiences of practitioners in using HPO to reach their goals, we asked the study participants to what extent they perceive to have attained which goals using what HPO methods.
Figure 7 shows the success rate per goal perceived by the survey participants. Roughly 75% of the participants responded to have successfully increased ML model performance, reached complied with target audience, increase ML model understanding, or satisfied requirements. Participants from industry were successful when trying to satisfy requirements or comply with target audience. 67% of the participants state that they were able to achieve the goal decrease practitioners effort. Only 62% of the participants perceived themselves as successful in decreasing necessary computations.
@@
The self-perceived success rates strongly vary between combinations of goals and HPO methods (see Figure 8). The survey participants perceived themselves as not successful when they tried to decrease necessary computations using manual tuning, grid search, and random search. They perceived themselves as rather successful in reaching this goal when using Bayesian optimization or evolutionary algorithms. Decreasing practitioner effort was best achieved using grid search or random search according to the survey participants. Bayesian optimization and evolutionary algorithms were perceived as less effective in decreasing effort. A potential explanation could be that those HPO methods often require more effort to be set up compared to others (Khuat et al., 2023). Moreover, participants also perceived manual tuning as ineffective in decreasing their efforts. Participants perceived manual tuning as very helpful to increase ML model understanding. Even though Bayesian optimization is considered a black-box optimization technique (Frazier, 2018), it was also perceived as suitable to increase ML model understanding. Random search, grid search, and evolutionary algorithms were perceived as unsuitable for increasing ML model understanding. Most participants did not perceive noteworthy differences between the effectiveness of HPO methods to successfully increase ML model performance and to satisfy requirements. Only evolutionary algorithms are significantly more successful when trying to satisfy requirements. Random search, grid search, and evolutionary algorithms were successfully used by survey participants to achieve comply with target audience. Bayesian optimization and manual tuning were applied with lower success rates for this goal.
@@ 5.1. Principal Findings
We conducted 20 semi-structured interviews and an online survey with 49 participants to better understand practitioner motives for using HPO methods. This study presents six principal goals (e.g.,  decrease practitioner effort, increase ML model understanding), which practitioners pursue in HPO of ML models, and fourteen contextual factors (e.g.,  HPO method comprehension, personal experience), which influence practitioners in their decisions for HPO methods to reach those goals.
@@
Notwithstanding the commonly high sample efficiency of Bayesian optimization, some practitioners prefer to use manual tuning to decrease the number of necessary computations. In particular, if practitioners assume that their ML model understanding is high, they expect to outperform Bayesian optimization. Yet, it is difficult to compare manual tuning to programmatic HPO methods due to its reliance on a mixture of explicit and implicit knowledge that often cannot be fully extracted from observations of practitioner actions. One way to leverage ML model understanding could be to integrate practitioner priors on the location of well-performing hyperparameter configurations into Bayesian Optimization effectively warm-starting optimizations (Souza et al., 2021; Hvarfner et al., 2022; Mallik et al., 2023).
Besides well-established goals pursued in HPO, this study presents a multitude of goals pursued by practitioners that are less emphasized in research on AutoML. Practitioners have strong motives for HPO beyond improving ML model performance. For example, the results of this study show that various practitioners are interested in better understanding their subject of work, including HPO methods and ML models, prior to using them. To better understand ML models, most practitioners chose HPO methods they understood over methods they would have to study first. This led practitioners to opt for manual tuning instead of programmatic HPO methods. Many participants perceived programmatic HPO tools as unsuited to increase ML model understanding. Even though many software packages for advanced HPO methods exist that can be used out of the box without understanding internals (e.g., SMAC3 (Lindauer et al., 2022) or Optuna (Akiba et al., 2019)), practitioners apparently are reluctant to use HPO methods they do not understand. Practitioners tend to rely on their own knowledge rather than giving up control to insufficiently understood ML models and HPO tools.
@@
The results presented in this study provide an aggregated view of practitioner motives in using HPO methods. Although our analysis did not reveal distinct personas with significantly different usage patterns, our results highlight nuanced variations in how practitioners engage with HPO methods. Specifically, we observed that the relevance of goals and the application of HPO methods remained relatively consistent across practitioners with varying levels of experience, age, and education. However, our findings indicate that contextual factors play a more prominent role in shaping the approaches of practitioners from academia versus those from industry (see Figure 9). Contrary to practitioners from industry, for example, academics tend to avoid using HPO methods that are difficult to integrate into workflows or require prior training. This especially applies when using complex HPO methods, according to the study participants.
@@
To make achievements of techno-centered research on HPO more actionable in practice and support development of novel and more user-centered HPO tools, it is important to incorporate practitioner motives for HPO into tool development. In response to calls for more human-centered AutoML (Lindauer et al., 2024; Lee and Macke, 2020; Sun et al., 2023; Xin et al., 2021), we propose improvements to programmatic HPO methods based on our findings to complement technological advances in AutoML, which primarily focus on conventional performance metrics, by incorporating social aspects.
@@ 5.3. Limitations and Future Work
The results presented in this work hint at possible conflicts between goals and contextual factors. Practitioners must resolve such conflicts to succeed in HPO, which entails prioritization of goals and the relevance of contextual factors. A possible conflict can arise when practitioners aim to ‘decrease necessary computations’ and ‘increase ML model performance’. Practitioners may attempt to find a Pareto-optimal achievement of both goals based on a clear prioritization. This work offers a foundation of goals and contextual factors that can may lead to trade-offs and call for prioritization of goals. Future work should investigate relationships between goals and contextual factors to uncover such trade-offs and investigate how practitioners resolve them (e.g., in terms of prioritization).
@@ 6. Conclusion
While programmatic HPO methods, such as Bayesian optimization and evolutionary algorithms, achieve high efficiency of the HPO process; practitioners sometimes opt for efficiency-wise inferior HPO methods, such as grid search and manual tuning. To understand practitioner motives for using HPO methods, we employed a two-step research approach consisting of semi-structured interviews and a survey based on an online questionnaire. We identified six principal goals pursued by practitioners in HPO, such as decrease practitioner effort, decrease necessary computations, and increase ML model understanding. Moreover, we extracted fourteen contextual factors that influence practitioners’ decisions for using HPO methods, such as available compute resources, HPO method traceability, and parallelization possibilities.
By bridging the gap between technological advancements and practitioner motives to use HPO methods, this work contributes to enhancing HPO practices and tools in the context of AutoML. In particular, the results of this study can be leveraged to guide the development of more user-centered HPO methods and HPO tools that cater to practitioner motives.

### v4

@@ 4.2. Perceived Success of Using HPO Methods to Achieve Specific Goals
Practitioners seem to have diverse and individual motivations for using specific HPO methods. However, these choices do not always lead to the desired outcomes. To distinguish between successful and unsuccessful experiences of practitioners in using HPO to reach their goals, we asked the study participants to what extent they perceive to have attained which goals using what HPO methods.
Figure 7 shows the success rate per goal perceived by the survey participants. Roughly 75% of the participants responded to have successfully increased ML model performance, reached complied with target audience, increase ML model understanding, or satisfied requirements. Industry participants were successful in trying to satisfy requirements or comply with target audience. 67% of the participants stated that they were able to achieve the goal decrease practitioners effort. Only 62% of participants considered themselves successful in reducing computational requirements.
@@
The self-perceived success rates strongly vary between combinations of goals and HPO methods (see Figure 8). Survey participants reported lower success rates in reducing computational requirements when using manual tuning, grid search, or random search. They perceived themselves as rather successful in reaching this goal when using Bayesian optimization or evolutionary algorithms. Decreasing practitioner effort was best achieved using grid search or random search according to the survey participants. Bayesian optimization and evolutionary algorithms were perceived as less effective in decreasing effort. A potential explanation could be that those HPO methods often require more effort to be set up compared to others (Khuat et al., 2023). Moreover, participants also perceived manual tuning as ineffective in decreasing their efforts. Participants perceived manual tuning as very helpful to increase ML model understanding. Even though Bayesian optimization is considered a black-box optimization technique (Frazier, 2018), it was also perceived as suitable to increase ML model understanding. Random search, grid search, and evolutionary algorithms were perceived as unsuitable for increasing ML model understanding. Most participants did not perceive noteworthy differences between the effectiveness of HPO methods to successfully increase ML model performance and to satisfy requirements. Only evolutionary algorithms were perceived as significantly more successful in trying to satisfy requirements. Survey participants successfully used grid search, random search, evolutionary algorithms, and Bayesian optimization to meet the expectations of their target audience. Manual tuning was applied with lower success rates for this goal.
@@ 5.1. Principal Findings
Notwithstanding the high sample efficiency of Bayesian optimization, some practitioners prefer to use manual tuning to decrease the number of necessary computations. In particular, if practitioners assume that their ML model understanding is high, they expect to outperform Bayesian optimization. Yet, it is difficult to compare manual tuning to programmatic HPO methods due to its reliance on a mixture of explicit and implicit knowledge that often cannot be fully extracted from observations of practitioner actions. One way to leverage ML model understanding could be to integrate practitioner priors on the location of well-performing hyperparameter configurations into Bayesian Optimization effectively warm-starting optimizations (Souza et al., 2021; Hvarfner et al., 2022; Mallik et al., 2023).
In addition to well-established goals pursued in HPO, this study presents a multitude of goals pursued by practitioners that are less emphasized in research on AutoML. Practitioners have strong motives for HPO beyond improving ML model performance. For example, the results of this study show that various practitioners are interested in better understanding their subject of work, including HPO methods and ML models, prior to using them. To better understand ML models, most practitioners chose HPO methods they understood over methods they would have to study first. This led practitioners to opt for manual tuning instead of programmatic HPO methods. Many participants perceived programmatic HPO tools as unsuited to increase ML model understanding. Although numerous software packages for advanced HPO methods are available for out-of-the-box use without requiring an understanding of their internal workings (e.g., SMAC3 (Lindauer et al., 2022) or Optuna (Akiba et al., 2019)), practitioners appear reluctant to adopt HPO methods they do not fully understand. Practitioners tend to rely on their own knowledge rather than giving up control to insufficiently understood ML models and HPO tools.
@@
The results presented in this study provide an aggregated view of practitioner motives in using HPO methods. Although our analysis did not reveal distinct personas with significantly different usage patterns, our results highlight nuanced variations in how practitioners engage with HPO methods. Specifically, we observed that the importance of goals and the use of HPO methods remained relatively consistent among practitioners, regardless of their experience, age, or education level. However, our findings indicate that contextual factors play a more prominent role in shaping the approaches of practitioners from academia versus those from industry (see Figure 9). Contrary to practitioners from industry, for example, academics tend to avoid using HPO methods that are difficult to integrate into workflows or require prior training. This especially applies when using complex HPO methods, according to the study participants.
@@
To make the achievements of techno-centric HPO research more actionable in practice and to support development of novel and more user-centered HPO tools, it is essential to integrate practitioners’ motivations for HPO into the tool development process. In response to calls for more human-centered AutoML (Lindauer et al., 2024; Lee and Macke, 2020; Sun et al., 2023; Xin et al., 2021), we propose improvements to programmatic HPO methods based on our findings to complement technological advances in AutoML, which primarily focus on conventional performance metrics, by incorporating social aspects.
@@ 5.3. Limitations and Future Work
The results presented in this work hint at possible conflicts between goals and contextual factors. Practitioners must resolve such conflicts to succeed in HPO, which entails prioritization of goals and the relevance of contextual factors. A possible conflict can arise when practitioners aim to ‘decrease necessary computations’ and ‘increase ML model performance’. Practitioners may attempt to find a Pareto-optimal achievement of both goals based on a clear prioritization. This work offers a foundation of goals and contextual factors that can lead to trade-offs and call for prioritization of goals. Future work should investigate relationships between goals and contextual factors to uncover such trade-offs and investigate how practitioners resolve them (e.g., in terms of prioritization).
@@ 6. Conclusion
While programmatic HPO methods, such as Bayesian optimization and evolutionary algorithms, achieve high efficiency of the HPO process; practitioners sometimes opt for efficiency-wise inferior HPO methods, such as grid search and manual tuning. To understand practitioner motives for using HPO methods, we performed a two-step research approach consisting of semi-structured interviews and a survey based on an online questionnaire. We identified six principal goals pursued by practitioners in HPO, such as decrease practitioner effort, decrease necessary computations, and increase ML model understanding. Moreover, we extracted fourteen contextual factors that influence practitioners’ decisions for using HPO methods, such as available compute resources, HPO method traceability, and parallelization possibilities.
By bridging the gap between technological advancements and practitioner motives to use HPO methods, this work contributes to enhancing HPO practices and tools in the context of AutoML. In particular, the results of this study can guide development of more user-centered HPO methods and HPO tools that cater to practitioner motives.
@@ Acknowledgements.
Appendix
Appendix A Interview Guide
The following presents the interview guide that we used in the semi-structured interviews.
A.1. Briefing
Dear study participant,
Thank you for supporting our research on hyperparameter optimization (HPO) in machine learning (ML) with your time and expertise. The interview will take about 30 minutes of your time. All data collected will be treated confidentially and reported only in aggregated form. Your responses will not be linked to your identity in any future publications.
HPO is an increasingly important topic in ML research, as it helps enhance model performance and enables objective comparisons of ML methods. In this study, we define HPO as the process of iteratively improving hyperparameter configurations during ML model training. A hyperparameter is a parameter whose value controls the learning process.
Various tools and ML libraries (e.g., Keras Tuner) support automated hyperparameter optimization. However, it remains unclear why practitioners often do not use advanced automated HPO [^3] Xavier Bouthillier, Gaël Varoquaux. Survey of machine learning experimental methods at NeurIPS2019 and ICLR2020. [Research Report] Inria Saclay Ile de France. 2020. https://hal.science/hal-02447823/document, and in which contexts manual HPO may even outperform automated optimization and vice versa. Our goal in this study is to understand the motives of practitioners for choosing different HPO methods.
All questions in this interview relate to your personal experience in ML, including the HPO methods you have used for hyperparameter optimization. To obtain detailed information on your work and its context, we ask you to refer to one ML model you developed for productive use in academia or industry. If possible, please provide a link to the publication or associated repository (e.g., GitHub).
The remainder of the interview is structured into three sections:
  1. (1)
    Hyperparameter Optimization in Machine Learning: We will discuss the ML project you will reference in your responses and seek to understand your choice of HPO methods.
  2. (2)
    Participant Background and Personal Experiences: We will ask about your demographics, background, and expertise in HPO and ML.
  3. (3)
    Debriefing: We will summarize the key findings from the interview and outline the next steps.
While we have prepared several questions, we welcome any additional insights you wish to share.
We appreciate your participation and encourage you to share your experiences and perspectives on HPO approaches. There are no wrong answers.
Thank you again for your time and participation!
A.2. Hyperparameter Optimization in Machine Learning
This section is divided into two subsections. First, we will ask you to select one of your previous ML projects to focus on during the interview and to describe the HPO methods you used. Second, we will ask about your reasons for choosing these methods.
A.2.1. Hyperparameter Optimization in a Selected Machine Learning Project
Details on Machine Learning Model
In the interview, we would like to refer to one of your previous ML projects. Please select an ML project that was meant for productive use in academia or industry, and where you optimized hyperparameters.
  1. (1)
    Optional: To which repository refer your following answers? Please provide the link to the repository of your model.
    URI:
  2. (2)
    In case your repository includes multiple models: please provide information to which exact ML model your following answers will refer.
  3. (3)
    Was you ML model empirically evaluated (e.g., in the form of a benchmark)?
    $\Box$ Yes  $\Box$ No
Optimization Approach
Next, we want to learn more about how you tuned hyperparameters of the selected ML model.
{mdframed}
[backgroundcolor=black!10] Interviewer Instruction: Question 4 must be answered ‘Yes’. Otherwise, the interviewee must not participate in the study.
  1. (4)
    Did you optimize the hyperparameters of the model selected for this interview?
    $\Box$ Yes  $\Box$ No
  2. (5)
    How did you tune your hyperparameters? If you have used a combination of different HPO methods, name all methods used.
    Exemplary answers are Bayesian optimization, grid search, manual tuning, and random search.
A.2.2. Motives to Use the Selected Hyperparameter Optimization Methods
{mdframed}
[backgroundcolor=black!10] Interviewer Instruction: This section builds on the responses from the previous section. The used HPO methods determine whether the questions need to be inverted by using ‘(not)’, as indicated in the following questions.
In the previous section, you indicated the HPO methods you used for your ML model: [LIST OF USED HPO METHODS]. We are now interested in your reasons for choosing these methods, as well as the advantages and disadvantages you experienced. We value your personal experiences and will not judge your responses—there are no wrong answers.
  1. (6)
    Which goals (e.g., increase ML model comprehension) did you try to reach through HPO by the chosen methods? A goal is an idea of the future or desired results where the achievement of the idea is decidable. If multiple HPO methods have been used, please explain in which order you used them and explain the individual goals you tried to achieve.
  2. (7)
    What potential advantages and disadvantages regarding the achievement of the described goals are you aware of with respect to the (combination of) HPO methods you have chosen? What potential advantages and disadvantages did you encounter in achieving these goals with the selected HPO methods? If you used multiple HPO methods, please specify which method each advantage and disadvantage refers to.
  3. (8)
    Why did you (not) choose random search? Please consider contextual factors (e.g., limited compute) and goals (e.g., improve model comprehension).
  4. (9)
    Why did you (not) manual tuning for HPO? Please consider contextual factors (e.g., limited compute) and goals (e.g., improve model comprehension).
  5. (10)
    Why did you (not) grid search for HPO? Please consider contextual factors (e.g., limited compute) and goals (e.g., improve model comprehension).
  6. (11)
    Why did you (not) choose Bayesian optimization for HPO? Please consider contextual factors (e.g., limited compute) and goals (e.g., improve model comprehension).
  7. (12)
    Which tools did you use for HPO? Exemplary tools are HPBandster, hyperopt, Ray, spearmint, and Ax.
  8. (13)
    How could future research support you in optimizing hyperparameters?
  9. (14)
    After you have helped us with your expertise, we would be grateful if you would share your thoughts on the contexts in which automated HPO may outperform manual HPO and vice versa.
A.3. Participant Background and Personal Experiences
In this section, we will ask you questions about your experience in artificial intelligence to better understand the context of your previous answers. Your personal information will not be disclosed in any way that could identify you.
  1. (15)
    What area(s) of artificial intelligence do you focus on?
    Exemplary answers are automatic speech recognition, computer vision, and natural language processing.
  2. (16)
    What is your main activity in artificial intelligence?
    Exemplary answers are consulting, development of statistical methods, library development, software solution development (i.e., using existing libraries and methods), and use case development.
  3. (17)
    What is/are your area(s) of expertise?
    Exemplary answers are AutoML, parallel computing, or (un-)supervised learning.
  4. (18)
    In which field do you work?
    Exemplary answers are academia, automotive, finance, IT support & services, and pharma.
  5. (19)
    How many years of experience do you have in ML?
    $\Box$ <2  $\Box$ 2-4  $\Box$ 5-7  $\Box$ 8-10  $\Box$ 11-15  $\Box$ >15
  6. (20)
    What is your highest educational level?
    Exemplary answers are Bachelor of Engineering, Bachelor of Science, Master of Science, and PhD.
  7. (21)
    What is your profession?
    Exemplary answers are big data engineer, developer, data scientist, and ML engineer.
  8. (22)
    What is your position in your organization?
    Exemplary answers are professor, PhD student, or lead software architect.
  9. (23)
    What is your age?
[…]
  10. (24)
    In which country are you primarily working or employed?
  11. (25)
    How many people are employed at your organization?
    $\Box$ <10  $\Box$ 10-50  $\Box$ 51-150  $\Box$ 150-500  $\Box$ 501-1000  $\Box$ 1,000
A.4. Debriefing
This is the final section of the interview. Thank you again for supporting our research with your valuable time and expertise.
{mdframed}
[backgroundcolor=black!10] Interviewer Instruction: Please summarize the key findings from the interview.
  1. (26)
    Would you like to receive a summary of the study results?
  2. (27)
    Do you know colleagues who might be interested in participating in this study? If yes, could you please connect us with them?
{mdframed}
[backgroundcolor=black!10] Interviewer Instruction: Outline the next steps.
We have now reached the end of the interview. Thank you very much for your participation and support!
Author Statement
This work is not under review at another journal or conference. It has not been published in conference proceedings or journal volumes. A previous version of this work was uploaded to a https://arxiv.org/abs/2203.01717.